OpenStudy (jessiegonzales):

MEDAL!!!! For questions 15 and 16, find the volume of the square pyramid shown. Round to the nearest whole number.

OpenStudy (jessiegonzales):

A.507cm^3 B.351cm^3 C.39cm^3 D.761cm^3

OpenStudy (jhannybean):

Well, Volume of a pyramid is \[V_P=(\text{base})^2\left(\frac{h}{3}\right)\]

OpenStudy (jessiegonzales):

vp=(13)^2(9/3)

OpenStudy (jhannybean):

Yes :)

OpenStudy (jhannybean):

The P was just to indicate it's the volume of the right pyramid, just a habit :P

OpenStudy (jessiegonzales):

vp=169(3)

OpenStudy (jhannybean):

And that equals?

OpenStudy (jessiegonzales):

A.507cm^3

OpenStudy (jhannybean):

good job :)

OpenStudy (jessiegonzales):

Mind helping with another?

OpenStudy (jhannybean):

Sure, close this question and open another :D